OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (mathstudent55):

x is the number of votes Maya received. y is the total number of votes cast

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

a total of 125,000 people voted: that means this is y

OpenStudy (mathstudent55):

Replace y with 125,000 and solve for x.

OpenStudy (anonymous):

yes so 0.25 x 125000 and then solve to get x=93750 @mathstudent55 and can u help me with another i think i did it correct i would just like reassurance

OpenStudy (mathstudent55):

x - 62,500 = 0.25y x - 62,500 = 0.25 * 125,000 x - 62,500 = 31,250 x = 93,750 You are correct.

OpenStudy (mathstudent55):

Yes. Pls start a new post for every new question.

OpenStudy (anonymous):

ok